if a population of rabbits consists of 9% white rabbits and 91% brown rabbits. Calculate all the genotypic frequencies of population A (size of 1, 000, 000) and population B (size of 1, 000)

Assumption: H-W conditions are held true
Answer #1

Answer:

Frequency of white=q^2= 9%=0.09

The frequency of white allele =q=sqrt of q^2= 0.3

The frequency of brown allele = p = 1-0.33=0.7 (as p+q=1)

Homozygous brown frequency=pp= 0.7*0.7=0.49

Heterozygous brown frequency=2pq=2*0.7*0.3=0.42

White frequency= q^2=0.09

Population A:

Homozygous brown = 0.49*1000000=490, 000

Heterozygous brown=0.42*1000000=420, 000

White = 0.09*1000000= 90, 000

Population B:

Homozygous brown= 0.49*1000=490

Heterozygous brown = 0.42*1000= 420

White= 0.09*1000= 90